A 215 g lead ball at a temperature of 83.1 ∘C is placed in a light calorimeter containing 158 g of water at 21.3 ∘C. Find the equilibrium temperature of the system.
Answer #1

Solution:

Given;

mass of lead m_{m}=215g

mass of water m_{w}=158g

Temperature of lead T_{m}=83.1^0C

Temperature of water T_{w}=21.3^0C

Specific heat capacity of lead c_{m}=0.0305cal/g^0C

Specific heat capacity of water c_{w}=1cal/g^0C

Let the final temperature of the system be T_{f} .

By the principle of calorimetery , we have;

  m_{m}c_{m}(T_{m}-T_{f}) =m_{w}c_{w}(T_{f}-T_{w})

  215\times0.0305\times(83.1-T_{f}) =158\times1\times(T_{f}-21.3)

  6.5575\times(83.1-T_{f}) =158\times(T_{f}-21.3)

544.92825-6.5575T_{f} =158T_{f}-3365.4

  164.5575T_{f} =3910.32825

T_{f} =\frac{3910.32825}{164.5575}=23.76^0C

The equilibrium temperature of the system is 23.76^0C .